# Promoting Constructive Deliberation: Reframing for Receptiveness

### v1

@@ 5.5 Toxicity effects on perceived reception
In this section, we analyze whether reframing may improve receptiveness at different rates based on the toxicity of the original reply. Toxicity scores $\in[0,1]$ are predicted using the Perspective API, which defines toxicity as “A rude, disrespectful, or unreasonable comment that is likely to make someone leave a discussion.” Prior studies have validated use of the Perspective API for toxicity classification on Reddit Rajadesingan et al. (2020). We use the mixed effects regression model as before but also include an interaction term for toxicity and strategy. As our previous regression did not reveal differences in receptiveness between strategies, we only compare differences between use of any strategy vs. baselines.
@@
We find that the difference in perceived receptiveness of our reframes based on the toxicity level of the original reply is statistically significant for some toxicity levels using the Ward test; the difference between low and high toxicity and between low and medium is significant, but the difference between medium and high toxicity is not. The differences between perceived receptiveness for the two baselines are not significant between any of the toxicity levels of the original replies. The mean differences between toxicity levels, test statistics, and p-values for both receptiveness reframes and baselines are shown in Table 13 in Appendix E.
These results suggest that the receptive reframes have higher impact on the more toxic comments. Interestingly, neither paraphrase nor receptive baseline generations were perceived to be any more receptive on toxic comments, further drawing the distinction between our reframing strategies and the baselines.
Figure 3: Receptiveness at different toxicity levels (per the PerspectiveAPI): low [0, 0.5), medium [0.5, 0.7), and high [0.7, 0.9]. Results are shown for the two baselines and the average over receptive-reframed replies. “B” indicates a baseline.
@@ 6 Discussion
As discussed in Section 5, our reframes have a smaller effect on reducing bias (F3) than on any of the other factors. This presents promising opportunities for future work on receptiveness reframing aimed at generating reframes that reduce outgroup bias. For example, this could include explicit instruction to reduce bias in addition to our prompts, or attempting to signal impartiality or humility AlSheddi (2020); Fisher (2000) as ways to implicitly reduce bias. However, these methods might result in a tradeoff between reducing bias and meaning preservation – discussed more in the limitations section.
@@
In this work, we only generate reframes using one strategy at a time, but future work could analyze combining multiple strategies. We do find that certain receptiveness reframes contain strategies that were not explicitly prompted for; this was especially prominent with hedges, which were commonly added to generations using other strategies. Understanding how different combinations of strategies might affect perceived receptiveness would allow us to refine the current framework. Future work could also automatically optimize for which strategies to use in different contexts.
@@
We have seen a vast increase in affective polarization, where increasingly, opposing partisans dislike and distrust each other Iyengar et al. (2019). This increase in affective polarization has contributed to partisan gridlock Hetherington and Rudolph (2020), and more recently, limited compliance with COVID-19 mitigation policies Druckman et al. (2021). A major intervention known to reduce polarization is actually fairly straightforward: to engage in conversations with the other side Levendusky and Stecula (2021). We believe that our computational framework for reframing, if integrated into conversation spaces, could improve the quality of conversations between opposing partisans, potentially reducing polarization Argyle et al. (2023).
@@
The ability to automatically reframe content could strengthen moderation at scale. Moderators could suggest receptiveness-signaling reframes to users as rewrites for comments which might be removed otherwise, discouraging users from commenting further (Myers West, 2018). Also, reframing one’s own comments provides an opportunity to learn from one’s mistakes, which has been shown to lead to better performance when learning a skill (Mason et al., 2016).
@@ 7 Conclusion
This work illustrates how transforming the social science construct of receptiveness into a computational framework can make LM generations more aligned with human perceptions. We found that our reframed replies are perceived to be better aligned with human notions of receptiveness than a reply generated with a generic receptiveness prompt. We also analyzed how our reframes are perceived in terms of the different receptiveness factors, and how some (emotion, curiosity, openness) are perceived to have improved more than others (bias). Finally, we demonstrated that our reframes appear to be more effective when the original replies are more toxic, and discuss applications of receptiveness-signaling for content moderation and reducing polarization.
Limitations & Ethical Considerations
There are several limitations and ethical considerations with this work.
@@
We curate six linguistic strategies from communications literature that signal receptiveness to another person. As we note previously, there is overlap between some of these strategies, particularly between agreement and grounding. However, given there are subtle differences in their implications, we decide to differentiate them. In addition, we choose these six strategies as a subset; there are potentially other strategies that we do not include here.
@@ Acknowledgements
We thank the online workers who provided annotations for this study. This research was supported in part by the Knight Foundation and by Good Systems [^5] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix A Prompts and Examples for Model Reframes
A.1 Prompts
@@ Acknowledgement
For the last strategy, we ultimately used “compromise and agreement” for <strategy>, instead of merely “agreement”, since the latter tended to cause the model to simply agree with the comment (even if this contradicted the original reply); adding “compromise” helped to remedy this. Each prompt is followed by the five in-context examples in A.2.
@@ A.2 Examples
Appendix B Additional Validation Results
We study differences in trigrams found in generated reframes compared to the original replies. The overlap of these trigrams between the different strategies is given in Table 5. Overlap is defined as whether a trigram in a strategy’s generations also appears in at least one other strategy’s generations.
Strategy Overlap
Hedging 0.318
Acknowledgement 0.340
Elaboration 0.283
Grounding 0.350
Gratitude 0.371
Agreement 0.364
Table 5: Proportion overlap between strategies.
Appendix C Additional Qualitative Examples
@@
Appendix D Receptiveness Factors
@@
Appendix E Statistics Tests
@@
Appendix F Annotation Interfaces

### v2

@@ 5.5 Toxicity effects on perceived reception
In this section, we analyze whether reframing may improve receptiveness at different rates based on the toxicity of the original reply. Toxicity scores $\in[0,1]$ are predicted using the Perspective API.
Prior studies have validated use of this API for toxicity classification on Reddit Rajadesingan et al. (2020). We use the mixed effects regression model as before but also include an interaction term for toxicity and strategy. As our previous regression did not reveal differences in receptiveness between strategies, we only compare differences between use of any strategy to the baselines.
@@
We find that the difference in perceived receptiveness of our reframes based on the toxicity level of the original reply is statistically significant (using the Ward test) for some toxicity levels; the difference between low and high toxicity and between low and medium is significant, but the difference between medium and high toxicity is not. The differences between perceived receptiveness for the two baselines are not significant between any of the toxicity levels of the original replies. The mean differences between toxicity levels, test statistics, and p-values for both receptiveness reframes and baselines are shown in Appendix Table 12.
These results suggest that the receptive reframes have higher impact on the more toxic comments. Neither paraphrase nor receptive baseline generations were perceived to be any more receptive on toxic comments, further drawing the distinction between our reframing strategies and the baselines.
Figure 3: Receptiveness at different toxicity levels (per the PerspectiveAPI): low [0, 0.5), medium [0.5, 0.7), and high [0.7, 0.9]. Results are shown for baselines B1 and B2 and the average over receptive-reframed replies.
@@ 6 Discussion
As discussed in Section 5, our reframes have a smaller effect on reducing bias (F3) than on any of the other factors. This presents promising opportunities for future work on receptiveness reframing aimed at generating reframes that reduce outgroup bias. For example, this could include explicit instruction to reduce bias in addition to our prompts, or attempting to signal impartiality or humility AlSheddi (2020); Fisher (2000) as ways to implicitly reduce bias. However, these methods might result in a tradeoff between reducing bias and meaning preservation. See Limitations for more discussion.
@@
While we only generate reframes using one strategy at a time, future work could explore combining strategies. We do find that certain receptiveness reframes contain strategies that were not explicitly prompted for; this was especially prominent with hedges, which were commonly added to generations using other strategies. Understanding how different combinations of strategies might affect perceived receptiveness would allow us to refine the current framework. Future work could also automatically optimize for which strategies to use in different contexts.
@@
Increasing affective polarization has contributed to partisan gridlock Hetherington and Rudolph (2020), and more recently, limited compliance with COVID-19 mitigation policies Druckman et al. (2021). A major intervention known to reduce polarization is actually fairly straightforward: to engage in conversations with the other side Levendusky and Stecula (2021). We believe that our computational framework for reframing, if integrated into conversation spaces, could improve the quality of conversations between opposing partisans, potentially reducing polarization Argyle et al. (2023).
@@
The ability to automatically reframe content could strengthen moderation at scale. Moderators could suggest receptiveness-signaling reframes to users as rewrites for comments which might be removed otherwise, discouraging users from commenting further (Myers West, 2018). Reframing one’s own comments also provides an opportunity to learn from mistakes, which is known to lead to greater improvement in skill learning (Mason et al., 2016).
@@ 7 Conclusion
This work illustrates how transforming the social science construct of receptiveness into a computational framework can make LLM generations more aligned with human perceptions. We found that our reframed replies are perceived to be better aligned with human notions of receptiveness than a reply generated with a generic receptiveness prompt. We also analyzed how our reframes are perceived in terms of the different receptiveness factors, and how some (emotion, curiosity, openness) are perceived to have improved more than others (bias). Finally, we demonstrated that our reframes appear to be more effective when the original replies are more toxic, and discuss applications of receptiveness-signaling for content moderation and reducing polarization.
Limitations
@@
We curate six linguistic strategies from social science literature that signal receptiveness to another person. There is overlap between some of these strategies, particularly agreement with acknowledgement and grounding. While the overlap with grounding was along expected lines (as discussed in Section 4), overlap with acknowledgement was unexpected. It appears that the agreement prompt does not always reframe replies with explicit agreement and instead appears to simply acknowledge the speaker’s perspective. This is likely because we prompt the model to make “explicit any agreement with the comment.” and in the absence of agreement expressed in the original reply, the model appears to default to acknowledge rather than to agree. For a more distinct strategy, we might instead study token agreement, where we don’t try to make any agreement explicit (because there usually isn’t any) and instead simply reframe in the form “yes, but…”.
@@ Acknowledgements
We thank the online workers who provided annotations for this study. This research was supported in part by the Knight Foundation and by Good Systems [^4] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. We acknowledge the Texas Advanced Computing Center (TACC) for compute resources. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
Appendix A Ensuring quality annotation
To ensure quality annotation, we apply two strategies to curate our annotator pool: (1) sourcing workers already pre-qualified by Amazon for its Ground Truth workforce [^5] https://aws.amazon.com/sagemaker/groundtruth/, and (2) filtering other workers ourselves, requiring at least 97% accuracy and 5000 completed hits. We also require workers to pass a qualification test consisting of four examples similar to our main task. Finally, given the US-centric nature of our data, we require workers to be US-based. We pay workers \$10/hr.
Appendix B Prompts and Examples for Model Reframes
B.1 Prompts
@@ Acknowledgement
For the last strategy, we ultimately used “compromise and agreement” for <strategy>, instead of merely “agreement”, since the latter tended to cause the model to simply agree with the comment (even if this contradicted the original reply); adding “compromise” helped to remedy this. Each prompt is followed by the five in-context examples in B.2.
@@ B.2 Examples
Appendix C Additional Validation Results
We study differences in trigrams found in generated reframes compared to the original replies. The overlap of these trigrams between the different strategies is given in Figure 4. Overlap between a pair of reframes is defined as the intersection of their trigrams divided by the union of their trigrams.
Figure 4: Pair-wise overlap between each of our strategy-reframes and baseline reframes. Overlap was calculated as the intersection of trigrams over the union, for each pair of reframes.
Appendix D Additional Qualitative Examples
@@
Appendix E Receptiveness Factors
@@
Appendix F Statistics Tests
@@
Appendix G Annotation Interfaces